Question 1 [20]

A simple gear train of 30 and 50 teeth for pinion and gear wheel respectively, is used to drive an electric wheelchair at a velocity that is double the maximum velocity of slip. The pinion has an equal addendum as the gear wheel, a module of 3 mm, and a pressure angle of 14.5o. The wheelchair has a maximum velocity of 5 m/s when its pinion rotates at 800 r.p.m. Calculate:

1.1. The

1.2. The

1.3. The

1.4. The

1.5. The

1.6. The

Question 2

magnitude of the velocity of slip of gears in mesh. (2) angular speed of the gear wheel in rad/s. (4) maximum length of a path of the approach in mm. (2) length of a path of a recess in mm. (7) addendum of the gears in mm. (7) contact ratio. (8)


1
Expert's answer
2021-10-10T09:32:13-0400

Solution;

Given;

No. of teeth of pinion,Tp=30

No. of teeth on gear,Tg=50

Module,m=3mm

Pressure angle,"\phi" =14.5°

Velocity of wheel chair ,vw=5m/s

Rotation of pinion,Np=800r.p.m

Pitch radius of pinion,r="\frac{mT_p}2" =45mm

Pitch radius of gear,R="\frac{mT_g}{2}=75mm"

(1.1) Magnitude of velocity of slip;

"v_s=2\u00d7v_w"

"v_s=2\u00d75m\/s" =10"m\/s"

(1.2) Angular speed of gear wheel;

From velocity ratio;

"\frac{N_p}{N_g}=\frac{T_g}{T_p}"

"N_g=\frac{N_p×T_p}{T_g}" ="\frac{800×30}{50}" =480rpm

Convert to angular velocity;

"w_g=\frac{2πN_g}{60}" ="\frac{2×π×480}{60}=50.27rad/s"

(1.3) maximum length of approach;

"P_a(max)=rsin\phi"

"P_a(max)=45sin(14.5\u00b0)"

"P_a(max)=11.27mm"

(1.4)path of recess in mm;

Addendum radius of pinion;

"r_a=r\sqrt{1+\frac Rr(\frac RR+2)sin^2\phi}"

"r_a=45\sqrt{1+\frac{75}{45}(\frac{75}{45}+2)sin^2(14.5°)}"

"r_a=52.92mm"

Hence,path of recess is ;

"p_r=\sqrt{r_a^2-r^2cos^2\phi}-rsin\phi"

"p_r=\sqrt{52.92^2-45cos^2(14.5)}-45sin(14.5)"

"p_r=18.7741mm"

(1.5) Addendum of the gears;

"A_g=R\sqrt{1+\frac rR(\frac rR+2)sin^2\phi}-1"

"A_g=75\sqrt{1+\frac{45}{75}(\frac{45}{75}+2)sin^2\phi}-1"

"A_g=3.585mm"

(1.6) contact ration,c.r;

"c.r=\frac{p_c}{C_p×cos\phi}"

pc is the path of contact

Cp is the circular pitch

Path of contact=path of recess +path of approach

"p_a=\sqrt{R_a-R^2cos^2\phi}-Rsin\phi"

"p_a=\sqrt{78.585^2-75^2cos^2(14.5)}-75sin(14.5)=11.2754mm"

"c.r=\frac{11.2754+18.7741}{π×3×cos(14.5°)}" =3.293